OpenStudy (anonymous):

Help. WILL FAN AND MEDAL. –1 + 5(6 – y) A. –y + 5 B. –y + 29 C. –5y + 29 D. –5y – 31

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@AutumnRoseT

OpenStudy (anonymous):

@badmood @briana.img

Vocaloid (vocaloid):

distribute the 5 and combine like terms

OpenStudy (anonymous):

?

Vocaloid (vocaloid):

5(6-y) = ?

OpenStudy (anonymous):

D?

Vocaloid (vocaloid):

not quite, just take it one step at a time what do you get when you simpify 5(6-y)? hint: a(b-c) = a*b - a*c

OpenStudy (anonymous):

30-y

Vocaloid (vocaloid):

not quite, you're very close remember to distribute the 5 to y

OpenStudy (anonymous):

30-5y??

Vocaloid (vocaloid):

good, so far we have -1 + 30 - 5y so, what do you get when you simplify?

OpenStudy (anonymous):

-y+29

Vocaloid (vocaloid):

not quite, remember the 5y doesn't change

OpenStudy (anonymous):

-5y+29

Vocaloid (vocaloid):

great! you found the answer :D

OpenStudy (anonymous):

I have more can you help me still?

Vocaloid (vocaloid):

sure, I guess

OpenStudy (anonymous):

ok I'll tag you.